Find the two numbers that if subtracted from 7 and 12,the product of the two numbers gives 104
Answer by ikleyn(53937)

.
Find the two numbers that if subtracted from 7 and 12,the product of the two numbers gives 104
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Your formulation is, OBVIOUSLY, incorrect.
Because if accept it LITERALLY, then the problem has INFINITELY many solution.


The correct formulation is THIS:

    Find the number which if subtracted from 7 and 12, the product of the two numbers gives 104


Then the problem is for MENTAL solution.

    Find two integer factors of the number 104 that multiply up 104 and differ by 12-7 = 5 units:


    104 = 2*52 = 4*26 = 8*13 . . .     Ooops ! You just found them: they are 8 and 13.


    Thus the unknown number under the question is (-1):


    7 - (-1) = 8;   12 - (-1) = 13   and  8*13 = 104.

Solved.